# Neural Audio Codecs for Prompt-Driven Universal Sound Separation

### v3

@@ 4.1 Results and Discussions
Sound Separation Performance on dnr-v2 (cf. Table 1). CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot) and retrained AudioSep+dnr-v2 across all categories, with sizable SI-SDR gains in speech (9.97 vs. 4.92/7.68 dB), music (1.15 vs. 2.46-2.46/5.55-5.55 dB), and SFX (0.89 vs. 0.34-0.34/4.66-4.66 dB). In ViSQOL, CodecSep matches or exceeds AudioSep in speech and music while slightly trailing on SFX, likely reflecting differences in SFX prompt distributions (AudioSep’s diverse training vs. CodecSep’s compositional SFX prompts from dnr-v2). Importantly, our bitstream-native variant—CodecSep+dnr-v2 (codes in: codes out, zero-shot)—evaluates the CodecSep+dnr-v2 masker directly on reconstructed embeddings $E$ from code streams (§3.1, para. 6677) without any fine-tuning: it incurs a modest drop relative to the continuous-latent path (about $1{\text{–}}2$ dB SI-SDR across sources; small ViSQOL deltas for music/speech and parity on SFX), yet still surpasses AudioSep+dnr-v2 on SI-SDR for all sources (e.g., music: 0.16-0.16 vs. 5.55-5.55 dB; speech: $8.26$ vs. $7.68$ dB; SFX: 0.96-0.96 vs. 4.66-4.66 dB). Compared to pretrained AudioSep (zero-shot), the codes-in:codes-out variant improves SI-SDR on speech/music but lags on SFX SI-SDR and ViSQOL. These results indicate that a deployment-friendly, no-finetuning bitstream path is already competitive; as discussed in §3.1, para. $6$ , light fine-tuning on $E$ or an embedding-consistency loss can close the residual gap. As part of our ablation, the lightweight CodecSep+dnr-v2 (ablate Masker) removes the transformer masker and applies FiLM directly to the encoder; it attains SI-SDR comparable to AudioSep+dnr-v2 with better perceptual speech quality, but overall separation quality drops due to FiLM perturbing the mixture latents held by the encoder.
@@
Effect of SFX Prompt Granularity during training. We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, CodecFormer, SDCodec) (cf. Table 2), (ii) generic 3-stem prompts (“music/speech/sfx”) (cf. Table 2), and (iii) universal prompting that retains generic cues for speech/music but uses fine-grained, compositional SFX descriptions (cf. Table 1). For (ii) and (iii), we train and evaluate separate versions of both CodecSep and AudioSep on each prompt setting. Across settings, CodecSep is competitive with or stronger than fixed-stem baselines on SFX while maintaining comparable speech/music quality (cf. Table 2). Under matched generic-prompt training/evaluation, CodecSep remains robust and surpasses spectrogram-domain AudioSep, indicating that effective separation does not hinge on carefully crafted prompts (cf. Table  1– 2). Moreover, replacing the “sfx” label with detailed SFX prompts consistently sharpens SFX extraction and, importantly, improves speech and music stems as well—suggesting that richer SFX supervision enhances overall scene disentanglement (cf. Table 1). While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
@@
TDANet SI-SDR $1.84^{\pm 3.55}$ $10.18^{\pm 2.91}$ $1.36^{\pm 4.90}$
ViSQOL $2.85^{\pm 0.58}$ $3.09^{\pm 0.43}$ $2.44^{\pm 0.72}$
CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
SDCodec SI-SDR $\mathbf{1.85^{\pm 3.68}}$ $\mathbf{11.32^{\pm 2.98}}$ $\mathbf{1.77^{\pm 4.08}}$
ViSQOL $\mathbf{2.96^{\pm 0.56}}$ $\mathbf{3.49^{\pm 0.40}}$ $\mathbf{2.64^{\pm 0.73}}$
@@
AudioSep SI-SDR $\mathbf{-2.46^{\pm 4.06}}$ $4.92^{\pm 4.21}$ $6.65^{\pm 4.73}$
(zero-shot) ViSQOL $\mathbf{2.86^{\pm 0.63}}$ $\mathbf{3.11^{\pm 0.56}}$ $2.08^{\pm 0.68}$
AudioSep + dnr-v2 SI-SDR $6.22^{\pm 2.77}$ $\mathbf{7.71^{\pm 3.11}}$ $2.11^{\pm 3.90}$
ViSQOL $2.55^{\pm 0.57}$ $2.47^{\pm 0.37}$ $2.39^{\pm 0.74}$
CodecSep + dnr-v2 SI-SDR $7.71^{\pm 2.84}$ $4.63^{\pm 2.48}$ $\mathbf{0.58^{\pm 4.15}}$
ViSQOL $2.45^{\pm 0.55}$ $2.70^{\pm 0.49}$ $\mathbf{2.39^{\pm 0.70}}$
@@
CodecFormer SI-SDR $5.67^{\pm 3.44}$ $2.27^{\pm 2.32}$ $6.54^{\pm 4.36}$
ViSQOL $2.16^{\pm 0.47}$ $2.51^{\pm 0.49}$ $2.13^{\pm 0.67}$
CodecSep + dnr-v2 SI-SDR $1.15^{\pm 3.35}$ $9.90^{\pm 2.91}$ $0.82^{\pm 4.18}$
(unguided, 3-stem) ViSQOL $2.75^{\pm 0.55}$ $3.09^{\pm 0.45}$ $\mathbf{2.48^{\pm 0.72}}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
(text-guided) ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
@@
AudioSep SI-SDR $7.75^{\pm 14.46}$ $8.57^{\pm 17.0}$ $7.62^{\pm 11.42}$ $7.03^{\pm 12.65}$ $6.44^{\pm 11.48}$
+ dnr-v2 ViSQOL $2.29^{\pm 1.12}$ $2.09^{\pm 1.08}$ $09^{\pm 1.00} 2.2$$22\pm 1.10$ $\mathbf{2.33^{\pm 1.08}}$
CodecSep SI-SDR $\mathbf{-5.87^{\pm 11.55}}$ $\mathbf{-6.02^{\pm 11.10}}$ $\mathbf{-6.37^{\pm 10.53}}$ $\mathbf{-6.12^{\pm 12.12}}$ $\mathbf{-6.09^{\pm 11.62}}$
+ dnr-v2 ViSQOL $\mathbf{2.34^{\pm 1.13}}$ $\mathbf{2.14^{\pm 1.09}}$ $\mathbf{2.15^{\pm 1.0}}$ $\mathbf{2.25^{\pm 1.11}}$ $2.24^{\pm 1.16}$
Robustness to prompt paraphrasing (cf. Table 5) . To probe lexical sensitivity, we re-evaluated both CodecSep + dnr-v2 and AudioSep + dnr-v2 on the dnr-v2 test split by replacing the generic training-time prompts for speech and music with three unseen paraphrases per class—speech: {“spoken voice”, “human conversation”, “people talking”}; music: {“instrumental music”, “band playing”, “melody with instruments”}. This constitutes a zero-shot paraphrase generalization test: models are trained with generic category cues but must respond to synonymic, potentially broader descriptors at inference. We observe a consistent qualitative pattern: (i) both systems exhibit the expected degradation when moving from generic to paraphrased prompts, confirming that lexical ambiguity weakens query–audio alignment; (ii) CodecSep degrades more gracefully overall, maintaining stronger separation and perceptual quality for speech, and retaining a small but persistent advantage for music; and (iii) the gap between models narrows under paraphrasing, yet the relative ranking is preserved, suggesting that FiLM-conditioned masking over structured codec latents confers robustness to synonym-level shifts. Notably, this study isolates lexical paraphrases; we did not incorporate paraphrases with explicit temporal qualifiers (e.g., “applause follows a song”), which we leave to future work.
Full inference complexity (cf. Table 6) . We compare end-to-end GMACs for spectrogram-domain separation (AudioSep) and our codec-latent approach (CodecSep). Code-stream input (typical deployment): when the server receives codec bitstreams, CodecSep runs a masker-only path and costs just $1.35$ GMACs, whereas AudioSep must decode, separate, and re-encode, totaling $73.6$ GMACs—about a $54\times$ compute reduction in favor of CodecSep. Architecture-only (separator compute, excluding codec): CodecSep remains at $1.35$ GMACs versus $33.5$ GMACs for AudioSep ($\approx 25\times$ lower), reflecting the smaller working representation in the codec latent space. Audio-stream input: if the server is fed raw audio, CodecSep incurs the codec overhead ($Enc(.)$ =12.28=12.28 GMACs, $Dec(.)$ =27.82=27.82 GMACs), yielding $41.45$ GMACs end-to-end versus $33.5$ GMACs for AudioSep; these codec costs disappear in the code-stream setting. Practically, even on the edge, full-band waveform processing is rarely feasible under memory constraints; audio is typically encoded to codes precisely to reduce footprint and bandwidth. Consequently, code-stream processing is the operational default, for which CodecSep is purpose-built.
Additional experiments (cf. Appendix DG) .
For readability, three extended studies are deferred to the Appendix, which also details data construction, prompt protocols (generic vs. universal and paraphrased variants), training/evaluation splits, and metrics. (i) Generalization across open-domain datasets (cf. Appendix D): On AudioCaps (derived from AudioSet), the spectrogram baseline AudioSep benefits from distributional alignment and attains the strongest absolute scores, yet CodecSep+dnr-v2 generalizes competitively—surpassing AudioSep+dnr-v2 in SI-SDR at comparable ViSQOL. When both are retrained on AudioCaps, CodecSep+AudioCaps again outperforms AudioSep+AudioCaps in separation quality. The same trend holds on the more challenging dnr-v2 test set, where mixtures often contain speech, music, and multiple overlapping SFX; both models drop in absolute performance, but CodecSep retains its advantage in SI-SDR. (ii) Relative–gain summaries (cf. Appendix E): We report percent improvements of CodecSep over AudioSep under matched training data and prompt settings; CodecSep shows large SI-SDR gains on dnr-v2 with positive (though smaller) gains under paraphrased prompts, and consistent improvements across ESC-50, Clotho-v2, AudioSet, VGGSound, and AudioCaps, with modest ViSQOL deltas (cf. Table 8). (iii) Higher bandwidth (cf. Appendix F): Swapping the frozen 16 kHz DAC for a 48 kHz (stereo–capable) EnCodec preserves the masking architecture and deployment path. Higher $F_{s}$ lengthens $T$ and adds high-frequency detail, making separation harder and lowering absolute SI-SDR/ViSQOL (cf. Table 9). Improvements hinge on backbone capacity/data—not on redesign. (iv) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided). We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original. This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix G (cf. Table 10).
Model Metric ($\uparrow$) Music Speech
AudioSep + dnr-v2 SI-SDR $6.43^{\pm 3.29}$ $4.14^{\pm 3.77}$
ViSQOL $2.53^{\pm 0.57}$ $2.64^{\pm 0.47}$
CodecSep + dnr-v2 SI-SDR $\mathbf{-5.60^{\pm 3.61}}$ $\mathbf{4.19^{\pm 4.18}}$
ViSQOL $\mathbf{2.56^{\pm 0.58}}$ $\mathbf{2.74^{\pm 0.51}}$
Table 6: Full Inference Complexity (GMACs $\downarrow$)
Model Input: Audio Stream Input: Code Stream Architecture-only
@@
CodecSep $41.45$ $\mathbf{1.35}$ ($\downarrow 54\times$) $\mathbf{1.35}$ ($\downarrow 25\times$)
Codec GMACs used above: DAC Enc(.) =12.28=12.28, Dec(.) =27.82=27.82.

### v4

@@ 4.1 Results and Discussions
Sound Separation Performance on dnr-v2 (cf. Table 1). CodecSep+dnr-v2 outperforms both pretrained AudioSep (zero-shot) and retrained AudioSep+dnr-v2 across all categories, with sizable SI-SDR gains in speech ($10.0$ vs. 4.94.9/$7.7$ dB), music (1.2 vs. 2.5-2.5/5.6-5.6 dB), and SFX ($0.9$ vs. 0.3-0.3/4.7-4.7 dB). In ViSQOL, CodecSep matches or exceeds AudioSep in speech and music while slightly trailing on SFX, likely reflecting differences in SFX prompt distributions (AudioSep’s diverse training vs. CodecSep’s compositional SFX prompts from dnr-v2). Importantly, our bitstream-native variant—CodecSep+dnr-v2 (codes in: codes out, zero-shot)—evaluates the CodecSep+dnr-v2 masker directly on reconstructed embeddings $E$ from code streams (§3.1, para. 6677) without any fine-tuning: it incurs a modest drop relative to the continuous-latent path (about $1{\text{–}}2$ dB SI-SDR across sources; small ViSQOL deltas for music/speech and parity on SFX), yet still surpasses AudioSep+dnr-v2 on SI-SDR for all sources (e.g., music: 0.2-0.2 vs. 5.6-5.6 dB; speech: $8.3$ vs. $7.7$ dB; SFX: 1.0-1.0 vs. 4.7-4.7 dB). Compared to pretrained AudioSep (zero-shot), the codes-in:codes-out variant improves SI-SDR on speech/music but lags on SFX SI-SDR and ViSQOL. These results indicate that a deployment-friendly, no-finetuning bitstream path is already competitive; as discussed in §3.1, para. $6$ , light fine-tuning on $E$ or an embedding-consistency loss can close the residual gap. As part of our ablation, the lightweight CodecSep+dnr-v2 (ablate Masker) removes the transformer masker and applies FiLM directly to the encoder; it attains SI-SDR comparable to AudioSep+dnr-v2 with better perceptual speech quality, but overall separation quality drops due to FiLM perturbing the mixture latents held by the encoder. Beyond AudioSep, CodecSep also outperforms the USS-pretrained BiModalSS model and the retrained text-conditioned Sudo rm-rf + FiLM baseline. The heavier attention-based conditioning used in BiModalSS does not generalize well to the open-domain mixtures in dnr-v2, leading to degraded performance. CodecSep also outperforms the Sudo rm-rf + FiLM model trained under the same universal setting. This gap is likely due to two factors. First, continuous CLAP embeddings provide far richer semantic conditioning than the fixed one-hot or multi-hot vectors that Sudo rm-rf + FiLM was designed for, causing the model to struggle with open-domain prompts. Second, applying FiLM across all U-Conv blocks while using only a single Conv1d layer for audio encoding can destabilize internal representations, especially when conditioned with high-dimensional continuous embeddings. Since AudioSep also outperforms both by a large margin, we use AudioSep as the primary baseline in our further experiments.
@@
Effect of SFX Prompt Granularity during training. We evaluate three regimes: (i) fixed-stem baselines without text guidance (TDANet, Sudo rm-rf, CodecFormer, SDCodec) (cf. Table 2), (ii) generic 3-stem prompts (“music/speech/sfx”) (cf. Table 2), and (iii) universal prompting that retains generic cues for speech/music but uses fine-grained, compositional SFX descriptions (cf. Table 1). For (ii) and (iii), we train and evaluate separate versions of both CodecSep and AudioSep on each prompt setting. Across settings, CodecSep matches or exceeds fixed-stem baselines on SFX while maintaining comparable speech and music quality, and its USS variant outperforms Sudo rm-rf across all stems with modest margins (cf. Table 2). Under matched generic-prompt training/evaluation, CodecSep remains robust and surpasses spectrogram-domain AudioSep, indicating that effective separation does not hinge on carefully crafted prompts (cf. Table  1– 2). Moreover, replacing the “sfx” label with detailed SFX prompts consistently sharpens SFX extraction and, importantly, improves speech and music stems as well—suggesting that richer SFX supervision enhances overall scene disentanglement (cf. Table 1). While these controlled studies cover multiple prompt granularities, we expect additional gains from training on larger, more diverse corpora with a spectrum of prompt specificities, which we leave for future work.
@@
TDANet SI-SDR $1.8^{\pm 3.55}$ $10.2^{\pm 2.91}$ $1.4^{\pm 4.90}$
ViSQOL $2.9^{\pm 0.58}$ $3.1^{\pm 0.43}$ $2.4^{\pm 0.72}$
Sudo rm-rf SI-SDR $0.9^{\pm 4.01}$ $9.0^{\pm 2.60}$ $0.6^{\pm 4.74}$
ViSQOL $2.7^{\pm 0.59}$ $2.9^{\pm 0.45}$ $2.3^{\pm 0.72}$
CodecFormer SI-SDR $5.7^{\pm 3.44}$ $2.3^{\pm 2.32}$ $6.5^{\pm 4.36}$
ViSQOL $2.2^{\pm 0.47}$ $2.5^{\pm 0.49}$ $2.1^{\pm 0.67}$
SDCodec SI-SDR $\mathbf{1.9^{\pm 3.68}}$ $\mathbf{11.3^{\pm 2.98}}$ $\mathbf{1.8^{\pm 4.08}}$
ViSQOL $\mathbf{3.0^{\pm 0.56}}$ $\mathbf{3.5^{\pm 0.40}}$ $\mathbf{2.6^{\pm 0.73}}$
@@
AudioSep SI-SDR $\mathbf{-2.5^{\pm 4.06}}$ $4.9^{\pm 4.21}$ $6.7^{\pm 4.73}$
(zero-shot) ViSQOL $\mathbf{2.9^{\pm 0.63}}$ $\mathbf{3.1^{\pm 0.56}}$ $2.1^{\pm 0.68}$
AudioSep + dnr-v2 SI-SDR $6.2^{\pm 2.77}$ $\mathbf{7.7^{\pm 3.11}}$ $2.1^{\pm 3.90}$
ViSQOL $2.6^{\pm 0.57}$ $2.5^{\pm 0.37}$ $2.4^{\pm 0.74}$
CodecSep + dnr-v2 SI-SDR $7.7^{\pm 2.84}$ $4.6^{\pm 2.48}$ $\mathbf{0.6^{\pm 4.15}}$
ViSQOL $2.5^{\pm 0.55}$ $2.7^{\pm 0.49}$ $\mathbf{2.4^{\pm 0.70}}$
@@
CodecFormer SI-SDR $5.8^{\pm 3.44}$ $2.3^{\pm 2.32}$ $6.5^{\pm 4.36}$
ViSQOL $2.2^{\pm 0.47}$ $2.5^{\pm 0.49}$ $2.1^{\pm 0.67}$
CodecSep + dnr-v2 SI-SDR $1.2^{\pm 3.35}$ $10.0^{\pm 2.91}$ $0.9^{\pm 4.18}$
(unguided, 3-stem) ViSQOL $2.8^{\pm 0.55}$ $3.1^{\pm 0.45}$ $\mathbf{2.5^{\pm 0.72}}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.2^{\pm 3.29}}$ $\mathbf{10.0^{\pm 2.92}}$ $\mathbf{0.9^{\pm 4.22}}$
(text-guided) ViSQOL $\mathbf{2.9^{\pm 0.57}}$ $\mathbf{3.2^{\pm 0.45}}$ $2.3^{\pm 0.73}$
@@
[…]
Additional experiments (cf. Appendix FJ) .
For readability, five extended studies are deferred to the Appendix, which also details data construction, prompt protocols (generic vs. universal and paraphrased variants), training/evaluation splits, and metrics. (i) Robustness to prompt paraphrasing (cf. Appendix F): We test lexical robustness by replacing the generic speech and music prompts with three unseen paraphrases each (e.g., “spoken voice,” “people talking”; “instrumental music,” “band playing”). This zero-shot paraphrase test requires models trained on generic labels to handle synonymic descriptors at inference. While both CodecSep+dnr-v2 and AudioSep+dnr-v2 degrade under paraphrasing, CodecSep remains consistently stronger (Table 6). This experiment focuses on lexical variants; paraphrases involving temporal structure (e.g., “applause follows a song”) are left for future work. (ii) Generalization across open-domain datasets (cf. Appendix G): On AudioCaps (derived from AudioSet), the spectrogram baseline AudioSep benefits from distributional alignment and attains the strongest absolute scores, yet CodecSep+dnr-v2 generalizes competitively—surpassing AudioSep+dnr-v2 in SI-SDR at comparable ViSQOL (cf. Table 7). When both are retrained on AudioCaps, CodecSep+AudioCaps again outperforms AudioSep+AudioCaps in separation quality. The same trend holds on the more challenging dnr-v2 test set, where mixtures often contain speech, music, and multiple overlapping SFX; both models drop in absolute performance, but CodecSep retains its advantage in SI-SDR. (iii) Relative–gain summaries (cf. Appendix H): We report percent improvements of CodecSep over AudioSep under matched training data and prompt settings; CodecSep shows large SI-SDR gains on dnr-v2 with positive (though smaller) gains under paraphrased prompts, and consistent improvements across ESC-50, Clotho-v2, AudioSet, VGGSound, and AudioCaps, with modest ViSQOL deltas (cf. Table 8). (iv) Higher bandwidth (cf. Appendix I): Replacing the frozen 16 kHz DAC with a 48 kHz, stereo-capable EnCodec leaves the masking architecture unchanged but increases latent-frame count and high-frequency detail, making separation harder and reducing SI-SDR/ViSQOL (Table 9). At higher $F_{s}$, gains depend mainly on backbone capacity and data rather than architectural modifications. Evaluations of 24 kHz and 44.1 kHz DAC variants show the same pattern: with fixed masker capacity, performance declines monotonically as sampling rate increases, reflecting the difficulty of wide-bandwidth separation, while DAC consistently outperforms EnCodec at matched bandwidths. (v) Reconstruction study: We probe source leakage with a single-source reconstruction diagnostic on dnr-v2 : each input contains one target source and models are either prompted with the matching caption (text-guided) or routed to the corresponding fixed head (non–text-guided). We also report mixture reconstruction by summing predicted source stems for a mixture and comparing to the original. This is a leakage/consistency check—not a primary separation metric; full setup/results are in Appendix J (cf. Table 10).
@@
AudioSep SI-SDR $7.8^{\pm 14.46}$ $8.6^{\pm 17.0}$ $7.6^{\pm 11.42}$ $7.0^{\pm 12.65}$ $6.4^{\pm 11.48}$
[…]
CodecSep SI-SDR $\mathbf{-5.9^{\pm 11.55}}$ $\mathbf{-6.0^{\pm 11.10}}$ $\mathbf{-6.4^{\pm 10.53}}$ $\mathbf{-6.1^{\pm 12.12}}$ $\mathbf{-6.1^{\pm 11.62}}$
+ dnr-v2 ViSQOL $\mathbf{2.3^{\pm 1.13}}$ $\mathbf{2.3^{\pm 1.09}}$ $\mathbf{2.2^{\pm 1.0}}$ $\mathbf{2.3^{\pm 1.11}}$ $2.2^{\pm 1.16}$
Table 5: Full Inference Compute Benchmarking Across Six Settings on 2424GB NVIDIA A-30 GPU
(a) Inference GMACs ($\downarrow$)
Model Audio Stream I/O Code Stream I/O Architecture-only
@@
Sudo rm-rf $\mathbf{16.44}$ $56.54$ $16.44$
CodecSep $41.45$ $\mathbf{1.35}$ $\mathbf{1.35}$
Codec GMACs: Enc=$12.28$, Dec=$27.82$
(b) Inference Time (s) ($\downarrow$)
Model Audio Stream I/O Code Stream I/O Architecture-only
AudioSep $\mathbf{0.33}$ $1.63$ $0.33$
Sudo rm-rf $0.51$ $1.81$ $0.51$
CodecSep $1.49$ $\mathbf{0.19}$ $\mathbf{0.19}$
Codec Inference Times (s): Enc=$1.16$, Dec=$0.14$
(c) Parameter count (in million) ($\downarrow$)
Model Audio Stream I/O Code Stream I/O Architecture-only
AudioSep $39$ $112.8$ $39$
Sudo rm-rf $\mathbf{5}$ $78.8$ $\mathbf{5}$
CodecSep $90.1$ $\mathbf{16.3}$ $16.3$
Codec Parameter Count (in million): Enc=$21.5$, Dec=$52.3$
(d) Parameter-only Memory Footprint (MB) ($\downarrow$)
Model Audio Stream I/O Code Stream I/O Architecture-only
AudioSep $156.13$ $447.62$ $156.13$
Sudo rm-rf $\mathbf{20.07}$ $311.25$ $\mathbf{20.07}$
CodecSep $339.89$ $\mathbf{48.4}$ $48.4$
Codec Parameter-only Memory Footprint (MB): Enc=$85.1$, Dec=$206.39$
(e) Forward/Backward Pass Memory Footprint (MB) ($\downarrow$)
Model Audio Stream I/O Code Stream I/O Architecture-only
AudioSep $\mathbf{804.3}$ $1580.6$ $804.3$
Sudo rm-rf $2032.13$ $2836.43$ $2032.13$
CodecSep $804.4$ $\mathbf{28.06}$ $\mathbf{28.06}$
Codec Forward/Backward Pass Memory Footprint (MB): Enc=$310.48$, Dec=$465.82$
(f) Full Memory Footprint (MB) ($\downarrow$)
Model Audio Stream I/O Code Stream I/O Architecture-only
AudioSep $\mathbf{960.43}$ $2028.22$ $960.43$
Sudo rm-rf $2052.2$ $3147.68$ $2052.2$
CodecSep $1144.25$ $\mathbf{76.46}$ $\mathbf{76.46}$
Codec Full Memory Footprint (MB): Enc=$395.58.$, Dec=$672.21$